OpenStudy (anonymous):

Find the product and simplify your answer. 9a^2(-5a^3 + 5a - 5)

OpenStudy (anonymous):

@amistre64 @BangkokGarrett @djcool31 @texaschic101 @mathmale

OpenStudy (anonymous):

@hoblos

OpenStudy (hoblos):

 in general a(b+c+d) = ab + ac + ad also, (a^x)(a^y) = a^(x+y) you must use these identities to solve your problem try solving it and tell me what you get, if you are still confused i'll help you through

OpenStudy (anonymous):

Yeah... I'm still confused lol... >.<

OpenStudy (hoblos):

lets do the first step 9a^2(-5a^3 + 5a - 5) = 9a^2(-5a^3) + 9a^2(5a) + 9a^2(-5) 9a^2(-5a^3) =? 9a^2(5a) + ? 9a^2(-5) = ? can you solve those ?

OpenStudy (anonymous):

Do you multiply for the first one? So it'd be... -45a^5?

OpenStudy (hoblos):

correct what about the other 2 ?

OpenStudy (anonymous):

Uhm.. -45a^5 45a^2 -45a^2 ?

OpenStudy (hoblos):

check the second one again

OpenStudy (anonymous):

OH! -45a^5 45a^3 -45a^2 ?

OpenStudy (hoblos):

now it's correct , so add them up and you'll get the final answer

OpenStudy (anonymous):

Ok.. uhh.. -45a^11 ?

OpenStudy (hoblos):

nope, we got 9a^2(-5a^3 + 5a - 5) = -45a^5 + 45a^3 -45a^2 so now we must find the common factor between the 3 terms do you know what a common factor is ?

OpenStudy (anonymous):

No I do not .-.

OpenStudy (hoblos):

the common factor is the factor found in all the terms example: xa + xb + xc the common factor is x so xa + xb + xc = x(a+b+c)

OpenStudy (anonymous):

Ok. So in this case the common factor is a?

OpenStudy (hoblos):

first do you have choices for the question to choose from ?

OpenStudy (anonymous):

No, you have to type it as a response =/.

OpenStudy (hoblos):

ok so you can just keep it in the simple form 9a^2(-5a^3 + 5a - 5) = -45a^5 + 45a^3 -45a^2

OpenStudy (anonymous):

So that's the final answer? O_o

OpenStudy (hoblos):

yes it's -45a^5 + 45a^3 -45a^2

OpenStudy (anonymous):

Thank you so much!! :)

OpenStudy (hoblos):

any time! :)